A mixture of 10 % acid and 90 % water is added to 5 liters of pure acid. The final mixture is 40 % water. How many liters of water are in the final mixture ?
Answer by stanbon(75887)

A mixture of 10 % acid and 90 % water is added to 5 liters of pure acid. The final mixture is 40 % water. How many liters of water are in the final mixture ?
--------------------
Equation:
acid + acid = acid
0.10x + 5 = 0.40(x+5)
10x + 500 = 40x + 200
30x = 300
x = 10 liters
--------------------------
Total liquid = x + 5 = 10+5 = 15 liters
Amount of water = 0.90x = 9 liters
--------------------------
Cheers,
Stan H.